My thirteen yr. old son is tired and having trouble with this Question
Find the number of degrees in the measure of an angle that is 20 degrees less than 4 times it's supplement. Thank You! John

Answer by mananth(16946)

let angle be x
supplement will be 180-x
...
x=4(180-x)-20
x=720-4x-20
x+4x=720-20
5x=700
x=140 deg the measure of the angle
..
m.ananth@hotmail.ca